• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Richard_Wilson TO

    Log entry time 23:04:35 on July16,2005

    Entry number 148256

    keyword=He target information

    The target was handed over at about 4.40 pm.  The He target only is being used 
    (loop 1.)
    
    It is cold with a temperature of 6.4 degrees all the intervening 6 1/2 hours.  
    
    The heat load is 0 from the beam (of course)
    14.31 from the fan 248 total
    
    
    
    


    FIGURE 1

    FIGURE 2